OpenStudy (anonymous):

Solve the equations and check your solutions. If there is no solution, say so ( I have the answer which is y=-1 I just to know the step by step please. thanks in advance) 2/3+5/(y-4)=(y+6)/(3y-12)

OpenStudy (anonymous):

\[\frac{2}{3}+\frac{5}{(y-4)}=\frac{(y+6)}{(3y-12)}\] rewrite as \[\frac{2}{3}+\frac{5}{(y-4)}=\frac{(y+6)}{3(y-4}\] multiply both side of the equation by \[3(y-4)\] in order to "clear" the fractions

OpenStudy (anonymous):

with the result \[2(y-4)+3(5)=y+6\]

OpenStudy (anonymous):

multiply and combine like terms with the result \[2y+7=y+6\] then finally isolating y we get \[y= -1\]

OpenStudy (anonymous):

no do not do it the hard way. there is no need to solve a quadratic equation here and in any case 4 is not a solution because you cannot replace y by 4 in the original expression

OpenStudy (anonymous):

exactly @Toriqur when you combine fractions the way you have, you introduce extra factors into the equation. In this case an extra factor of y-4 the least common denominator for the entire equation is 3(y-4)

OpenStudy (anonymous):

ebbflo (nice) has it

OpenStudy (anonymous):

satellite73: thanks for chiming in, this thread is more or less dead but its good to have good information none the less. I have witnessed teachers instructing their students to combine fractions in exactly that manner because(i am guessing) it works out fine in the end when adding numerical fractions.

OpenStudy (anonymous):

math teachers generally suck. if i see another question asking to "simplify" i am going to chuck a fit